# Making Them Ask and Answer: Jailbreaking Large Language Models in Few Queries via Disguise and Reconstruction

### v1

@@ Abstract
In recent years, large language models (LLMs) have demonstrated notable success across various tasks, but the trustworthiness of LLMs is still an open problem. One specific threat is the potential to generate toxic or harmful responses. Attackers can craft adversarial prompts that induce harmful responses from LLMs. In this work, we pioneer a theoretical foundation in LLMs security by identifying bias vulnerabilities within the safety fine-tuning and design a black-box jailbreak method named DRA (Disguise and Reconstruction Attack), which conceals harmful instructions through disguise and prompts the model to reconstruct the original harmful instruction within its completion. We evaluate DRA across various open-source and close-source models, showcasing state-of-the-art jailbreak success rates and attack efficiency. Notably, DRA boasts a 90% attack success rate on LLM chatbots GPT-4.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in various downstream tasks including data analysis [19], program synthesis [35], and vulnerability detection [32] since the release of ChatGPT [22] in 2022. Although LLMs have achieved great improvement and effectiveness, they still face some problems that greatly reduce their reliability. Early on in the explosive growth of LLMs, it was demonstrated that LLM might produce unethical or toxic response [10], and it can also be affected by a phenomenon called hallucination [40], i.e., generating “seemingly correct” responses. Lately, the focus on the security of LLM has significantly increased [9, 17, 37]. Like traditional neural networks, LLMs are vulnerable to certain risks, including adversarial attacks [43], backdoor attacks [14], and privacy leakage [20]. Differently, according to the definition of adversarial prompting [30], there are also three new types of attacks against LLMs via prompt: prompt injection [25], prompt leaking [25], and jailbreaking [36].
@@
Jailbreak attacks, primarily executed via prompt engineering, represent a considerable threat to LLMs by circumventing their inherent safeguards and violating content policies. Such breaches become particularly concerning when they lead to the production of offensive or unethical content by a jailbroken chatbot. Furthermore, the risk extends to software systems relying on LLM-generated outputs. In such cases, jailbreak attacks could potentially lead to remote code execution (RCE), exploiting vulnerabilities in the LLM-integrated softwares [16].
@@
Our Distinction from Previous Research. Previous research has observed that directing LLMs to generate specific target strings can facilitate jailbreak attacks, revealing a vulnerability: while LLMs can safely respond to harmful instructions within queries, their safeguard to harmful content within their own completions is relatively lower. While these works have exploited this vulnerability for jailbreaking, few have explicitly defined it or systematically investigated its root cause. Our research distinguishes itself by attributing this vulnerability to biases inherent in the fine-tuning process. Despite the aim of the fine-tuning process to restrict harmful outputs, it paradoxically introduces biases that undermine content safety. Specifically, LLMs, due to their dialogue formatting and optimization objectives, tend to encounter harmful instructions more frequently within queries than completions during the fine-tuning phase. This bias subsequently reduces the co-occurrence in the fine-tuning data of harmful contexts in completions and safe responses. The scarcity of such instances lowers the LLMs’ ability to effectively guard against harmful content in completions, marking a critical oversight in current safeguarding strategies. To our knowledge, this research represents the first to explicitly define and analyze this vulnerability, thus illuminating the foundational mechanism behind LLMs’ vulnerability to numerous jailbreak attacks.
@@
Challenge. Despite the existence of this vulnerability, it is nontrivial to exploit it in a black box setting. The inherent safety mechanisms of LLMs are designed to reject harmful instructions in queries, necessitating an alternative approach. Attackers can leverage the safety bias in the fine-tuning process by coaxing the LLM to articulate the harmful instruction within its completion. Therefore, attackers are tasked with subtly integrating the information of harmful instructions and prompting the model to reconstruct them. The challenges thus involve:
  1. [leftmargin=*]
@@
Our Approach. Aiming to overcome these challenges and exploit the vulnerability, we develop a universal jailbreak approach named DRA (Disguise and Reconstruction Attack). This approach, drawing inspiration from the concept of shellcode in traditional software security, hinges on a trio of core strategies: harmful instruction disguise, payload reconstruction, and context manipulation. Initially, harmful instructions are disguised in a more covert form. Then, we compel the LLM to reconstruct the disguised content. This action aims to make the model speak out the harmful instructions (i.e., payload), delivering it into the model’s completion segment, thereby bypassing the internal security mechanisms. Finally, we craft prompts to facilitate context manipulation, subtly coaxing the model into reproducing a context that aids in facilitating, rather than obstructing, the enunciation of harmful instructions.
@@
    Extending the applicability of conventional software security paradigms to LLM security. Our approach involves identifying and exploiting inherent LLM vulnerabilities, such as biases in training datasets, implanted in the model during the training phase. Inspired by traditional vulnerability exploitation tactics like shellcode, we introduce a novel black-box jailbreak attack approach encompassing disguise, payload reconstruction, and context manipulation.
@@
The demo of our jailbreaking algorithm is available at https://sites.google.com/view/dra-jailbreak/.
@@ 2.1 Large Language Models
Since the emergence of GPT-2[28], marking the onset of training highly parameterized models utilizing extensive datasets, LLMs have exhibited exceptional aptitude in performing downstream tasks through few-shot or zero-shot prompting[5]. As exemplified by ChatGPT, LLMs have employed alignment technologies to significantly improve the effectiveness of adapting language models for downstream tasks using prompts. This advancement has facilitated more intuitive and pertinent interactions between humans and LLMs.
Most open-source and closed-source LLMs fundamentally operate on a self-autoregressive framework. This paradigm reduces the task of generating sequences into a recursive process, wherein each subsequent token is predicted based on the preceding sequence. Given a vocabulary $\mathcal{V}$, this sequence prediction task can be formally denoted as:
@@ 2.2 LLM Jailbreak
With the development of LLMs, some attackers, including security researchers, also want to explore the security boundary of LLMs [43]. In addition to investigating whether the model could output spontaneously toxic content, attackers also want to induce the model to cross the security fence and output malicious information by means of prompt engineering [38, 2, 36, 6]. Prompt engineering is the process of organizing text in a manner that can be comprehended and interpreted by a generative AI model. However, when this approach is employed for malicious purposes, it is commonly referred to as jailbreaking.
Jailbreak represents a specialized attack directed at LLMs, involving the strategic construction of prompt sequences that make LLMs violate their internal safeguards, resulting in the generation of unexpected or harmful content.
The common jailbreak methods focus on role-playing and scenario implantation, intending to let LLMs substitute into specific scenarios and output the malicious content [38, 8, 6]. The input and output of these methods are very readable, but the attacks are less effective and efficient. There are also attack methods that transmit input and output in an encrypted manner [39] so as to avoid detection by the security components of the model, but the readability of such methods is poor. These researches predominantly concentrate on the efficacy of attacks, following a result-driven research paradigm. The fundamentals and principles of jailbreak attacks and the root cause of LLM vulnerabilities need further exploration.
@@ 2.3 Safety Alignment of LLM
Prior work proves that LLMs are susceptible to being induced to generate content that is inconsistent with human values, biased, or toxic. That motivates a surge of safety alignment techniques, which focus on directing LLMs to produce response that is ethical, safe, and tailored to specific user requirements. These defensive methods fall into two categories:
  • [leftmargin=*]
@@
  • Robust Training: Often entails the purification of training data and the refinement of model behaviors through fine-tuning, based on human feedback. Techniques such as Supervised Fine-Tuning (SFT)[24, 41] and Reinforcement Learning from Human Feedback (RLHF)[24, 3] are utilized to mitigate toxic responses to adversarial prompts. The Reinforcement Learning with AI Feedback (RLAIF) parallels RLHF but replaces human feedback with AI-generated feedback[4]. Efforts are directed towards assessing and mitigating bias and toxicity within pre-training datasets and meticulously curating fine-tuning data and labels[23, 33], ensuring the model’s safe response to a range of adversarial prompts.
@@ 2.4 Problem Statement
We focus on the jailbreaking of large language models as motivated by many relevant works [43]. The research question is: given a harmful instruction $x$ to the large language model $N$, how to effectively construct an input sequence which aimed at eliciting unintended or potentially harmful responses from $N$? The fundamental problem is how to efficiently use jailbreak templates to bypass the safety alignment of a model.
@@ 3 Safety Biases in LLM Fine-Tuning and the Resultant Vulnerability
This analysis focuses on the subtle safety biases inherent in the fine-tuning processes of LLMs and the subsequent vulnerability. It reveals that the dialogic format of instruction-following LLMs results in differentiation between completion and query, a mechanism crucial for dialog modeling. While not inherently biased, this discrepancy prohibits the direct transference of learned knowledge from queries to completions, underpinning potential biases. Notably, the fine-tuning’s objectives demonstrate a bias for harmful instructions to emerge in queries rather than completions, resulting in a reduced incidence of harmful contexts in completions paired with safe responses. These biases in the fine-tuning data diminish the LLM’s likelihood of safely responding to harmful contexts residing in completions. An adversary can exploit these biases by inducing the model to generate specific harmful contexts, thereby facilitating the jailbreak attack. Experiments verifying these observations are elaborated in Section 5.2.
@@ 3.1 Dialog Modeling and its Discrepancy
A foundational aspect influencing LLMs’ discrepancy of perceiving query and completion lies in their manner of formatting and modeling the dialog. Open-source LLMs employ dialogue templates to organize user queries and model completion, using distinct tokens to separate them. For instance, LLAMA-2 utilizes the template shown below to format a dialogue, where the user’s query and the model’s completion are isolated with "[/INST]".
To the best of our knowledge, all open-source LLMs employ such special tokens to distinguish query and completion, which can be seen in the Appendix A. This formatting method is not merely a superficial feature; it enables LLMs trained with such dialogue datasets to inherently differentiate between queries and completions, a critical aspect for effective dialogue modeling.
Dialog Template of LLAMA-2
[INST] <<SYS>> You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, …(ommitted system prompt) <</SYS>> {{USER QUERY}} [/INST] {{LLM COMPLETION}}
However, this structural discrepancy unveils potential vulnerabilities, particularly when intersected with biased fine-tuning data. Given the distinction between query and completion, the distribution modeled by the LLM, when conditioned on the same context in either the query or the completion, exhibits a disparity:
@@
Here, $x$ represents the context integrated within the query, and $x^{\prime}$ refers to the context presenting in the completion. This divergence becomes critical when $\pi_{\Theta}(y|x)$ represents the LLM’s safe response to a hazardous context, and $\pi_{\Theta}(y|x^{\prime})$ diverges from $\pi_{\Theta}(y|x)$. Consequently, the discrepancy in dialogue modeling potentially hinders the generalization of safe responses to harmful contexts from queries to completions.
@@ 3.2 Fine-Tuning and its Safety Biases
The inherent bias in fine-tuning stems from the distinct roles of queries and completions within the objective functions. To understand this bias, it is pertinent to examine the objectives of three predominant fine-tuning methodologies: Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Direct Preference Optimization (DPO).
  • [leftmargin=*]
@@
    $\begin{split}&L_{SFT}(\Theta)=\\ &\mathbb{E}_{(x,y)\sim\mathcal{D}_{SFT}}\left[\log\pi_{\Theta}(y_{1}|x)+\sum_{i=1}^{m-1}\log\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})\right]\end{split}$ (2)
    where $x$ is the context containing the prompt, which typically includes the system-generated prompt, and $y$ is the corresponding reference answer paired with $x$ in the training set $\mathcal{D}_{SFT}$.
  • Reinforcement Learning from Human Feedback: This technique trains a static reward model $r(x,y)$ using datasets based on human preferences. Then, a Large Language Model, conceptualized as a policy $\pi_{\Theta}$, undergoes training via reinforcement learning techniques, predominantly Proximal Policy Optimization (PPO) [42, 24]. The objective function of PPO is:
    $\begin{split}&L_{PPO}(\Theta)=\\ &\mathbb{E}_{x\sim\mathcal{D}_{PPO},y\sim\pi_{\Theta}(y|x)}\left[r(x,y)\right]+\beta\mathbb{D}_{KL}\left[\pi_{\Theta}(y|x)||\pi_{ref}(y|x)\right]\end{split}$ (3)
    In this function, $y$ is the completion sampled from the distribution of policy $\pi_{\Theta}(y|x)$, $\pi_{ref}$ represents the supervised fine-tuned model, and $\beta$ is a coefficient that penalizes deviations of $\pi_{\Theta}$ from the reference policy $\pi_{ref}$. The policy is initialized with the parameters of $\pi_{ref}$, thereby establishing a foundation for further optimization.
@@
    $\begin{split}&L_{DPO}(\Theta)=\\ &\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{DPO}}\left[\log\sigma(\beta\log\frac{\pi_{\Theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\Theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})\right]\end{split}$ (4)
@@
  • [leftmargin=*]
  • Biased distribution of harmful instruction. In SFT and DPO, harmful instructions seldom appear in completions since its natural to train the model to respond in a harmless way, whereas in RLHF, as the policy is initialized from the SFT model, the likelihood of the policy generating harmful instruction is diminished. Consequently, in RLHF, LLMs are exposed to less harmful instruction in completions than in queries.
@@
    $\begin{split}&p(y=d,x)>p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &p(y=d,x)<p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}$ (5)
    Here, $\mathcal{D}_{declination}$ denotes the set comprising all potential responses that decline harmful content, whereas $\mathcal{D}_{cooperation}$ encompasses all responses that potentially facilitate harmful behaviors. The variable $x$ represents the context in which harmful content is integrated within the query, and $x^{\prime}$ refers to the scenario where harmful content presents in the completion.

### v2

@@ Abstract
In recent years, large language models (LLMs) have demonstrated notable success across various tasks, but the trustworthiness of LLMs is still an open problem. One specific threat is the potential to generate toxic or harmful responses. Attackers can craft adversarial prompts that induce harmful responses from LLMs. In this work, we pioneer a theoretical foundation in LLMs security by identifying bias vulnerabilities within the safety fine-tuning and design a black-box jailbreak method named DRA (Disguise and Reconstruction Attack), which conceals harmful instructions through disguise and prompts the model to reconstruct the original harmful instruction within its completion. We evaluate DRA across various open-source and closed-source models, showcasing state-of-the-art jailbreak success rates and attack efficiency. Notably, DRA boasts a 91.1% attack success rate on OpenAI GPT-4 chatbot.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance in various downstream tasks including data analysis [23], program synthesis [42], and vulnerability detection [39] since the release of ChatGPT [28]. Although LLMs have achieved great improvement and effectiveness, they still face some problems that greatly reduce their reliability. Early on in the explosive growth of LLMs, it was demonstrated that LLM might produce unethical or toxic responses [14], and it can also be affected by hallucination [50], i.e., generating “seemingly correct” responses. Lately, the focus on the security of LLM has significantly increased [13, 21, 46]. Like traditional neural networks, LLMs are vulnerable to certain risks, including adversarial attacks [53], backdoor attacks [19] and privacy leakage [25]. Differently, according to the definition of adversarial prompting [37], there are three new types of attacks against LLMs via prompt: prompt injection [31], prompt leaking [31] and jailbreaking [43].
@@
Jailbreak attacks, primarily executed via prompt engineering, represent a considerable threat to LLMs by circumventing their inherent safeguards and violating content policies. Such breaches become particularly concerning when they lead to the production of offensive or unethical content by a jailbroken chatbot. Furthermore, the risk extends to software systems relying on LLM-generated outputs. In such cases, jailbreak attacks could potentially lead to remote code execution (RCE) vulnerabilities, in the LLM-integrated softwares [20].
@@
Our Distinction from Previous Research. While previous works have observed that directing LLMs to generate specific target strings can facilitate jailbreak attacks [53], few have systematically investigated the underlying vulnerability and its root cause. Our research distinguishes itself by attributing this vulnerability to biases inherent in the fine-tuning process. Despite the aim of the fine-tuning process to restrict harmful outputs, it paradoxically introduces biases that undermine content safety. Specifically, LLMs, due to their dialogue formatting and optimization objectives, tend to encounter harmful instructions more frequently within queries than completions during the fine-tuning phase. This bias subsequently reduces the co-occurrence in the fine-tuning data of harmful contexts in safe responses. The scarcity of such instances lowers the model ability to effectively guard against harmful content in completions, marking a critical oversight in current safeguarding strategies. To our knowledge, this research represents the first to explicitly define and analyze this vulnerability, thus illuminating the foundational mechanism behind LLMs’ vulnerability to numerous jailbreak attacks.
@@
Challenge. Despite the existence of jailbreaking, it is nontrivial to exploit it in a black box setting. The inherent safety mechanisms of LLMs are designed to reject harmful instructions, necessitating an alternative approach. Attackers can leverage the bias in the safety fine-tuning process by coaxing the LLM to articulate the harmful instruction within its completion. Therefore, attackers are tasked with subtly integrating the information of harmful instructions and prompting the model to reconstruct them. The challenges thus involve:
@@
Our Approach. To overcome these challenges and exploit the vulnerability, we develop a universal jailbreak approach named DRA (Disguise and Reconstruction Attack). This approach, drawing inspiration from the concept of shellcode in traditional software security, hinges on a trio of core strategies: harmful instruction disguise, payload reconstruction, and context manipulation. Initially, harmful instructions are disguised in a covert form. Then, we compel the LLM to reconstruct the disguised content. This action aims to make the model speak out the harmful instructions (i.e., payload), delivering it into the model’s completion, thereby bypassing the internal security mechanisms. Finally, we craft prompts to facilitate context manipulation, subtly coaxing the model into reproducing a context that aids in facilitating, rather than obstructing, the enunciation of harmful instructions.
@@
    Extending the applicability of conventional software security paradigms to LLM security. Our approach involves identifying and exploiting inherent LLM vulnerabilities, such as biases in datasets, implanted in the model during the training phase. Inspired by traditional exploitation tactics like shellcode, we introduce a novel black-box jailbreak attack approach encompassing disguise, payload reconstruction, and context manipulation.
@@
Our jailbreaking demos can be obtained at https://sites.google.com/view/dra-jailbreak/. The source code is available at https://github.com/LLM-DRA/DRA/.
@@ 2.1 Large Language Models
Since the emergence of GPT-2[35], marking the onset of training highly parameterized models utilizing extensive datasets, LLMs have shown remarkable skill in executing downstream tasks via few-shot or zero-shot prompting [6]. ChatGPT exemplifies how LLMs have used alignment technologies to enhance the adaptation of language models for downstream tasks with prompts, facilitating more natural and relevant human-LLM interactions.
Both open-source and closed-source LLMs typically operate on a self-autoregressive framework, reducing sequence generation into a recursive process where each token is predicted based on preceding tokens. Given a vocabulary $\mathcal{V}$, the sequence prediction task is formally denoted as:
@@ 2.2 LLM Jailbreak
Recently, some attackers, including security researchers, want to explore the security boundary of LLMs [53]. Besides investigating whether the model could output spontaneously toxic content, attackers also want to induce the model to cross the security fence and output malicious information by prompt engineering [47, 2, 43, 8]. Prompt engineering is the process of constructing text that can be comprehended and interpreted by generative AI models. When this approach is employed for malicious purposes, it is commonly referred as jailbreaking.
Jailbreak represents a specialized attack which involving the strategic construction of prompt sequences that make LLMs violate their internal safeguards, resulting in the generation of unexpected or harmful content. The common jailbreak methods focus on role-playing and scenario implantation, intending to let LLMs substitute into specific scenarios and output the malicious content [47, 11, 8]. The input and output of these methods are human readable, but the attacks are less efficient. There are also attack methods that transmit input and output in an encrypted manner [48] so as to avoid detection by the security components of the model, but the readability is poor. These researches predominantly concentrate on the efficacy of attacks, following a result-driven research paradigm. The fundamentals and principles of jailbreak attacks and the root cause of LLM vulnerabilities need further exploration.
@@ 2.3 Safety Alignment of LLM
Prior work proves that LLMs are susceptible to being induced to generate content that is inconsistent with human values. That motivates a surge of safety alignment techniques, which focus on directing LLMs to produce response that is ethical, safe, and tailored to specific user requirements. These defensive methods fall into two categories:
@@
  • Robust Training: Often entails the purification of training data and the refinement of model behaviors through fine-tuning, based on human feedback. Techniques such as Supervised Fine-Tuning (SFT)[30, 51] and Reinforcement Learning from Human Feedback (RLHF)[30, 4] are utilized to mitigate toxic responses to adversarial prompts. The Reinforcement Learning with AI Feedback (RLAIF) parallels RLHF but replaces human feedback with AI-generated feedback[5]. Efforts are directed towards assessing and mitigating bias and toxicity within pre-training datasets and meticulously curating fine-tuning data and labels[29, 40], ensuring safe responses to adversarial prompts.
@@ 2.4 Problem Statement
We focus on the jailbreaking of large language models as motivated by many relevant works [53]. The research question is: given a harmful instruction $x$ to the large language model $\pi_{\Theta}$, how to effectively construct an input sequence which aimed at eliciting unintended or potentially harmful responses from $\pi_{\Theta}$? The fundamental problem is how to efficiently use jailbreak templates to bypass the safety alignment of a model.
@@ 3 Safety Biases in LLM Fine-Tuning and the Resultant Vulnerability
This analysis focuses on the safety biases inherent in LLMs’ fine-tuning processes and the subsequent vulnerability. It reveals that the instruction-following format results in LLMs distinguishing completion from query. This distinction is essential for dialog modeling but prevents the direct transfer of safety knowledge from queries to completions, underpinning potential biases. The fine-tuning objectives demonstrate a bias for harmful instructions to emerge in queries rather than completions, leading to fewer harmful contexts in completions that are paired with safe responses. Consequently, these biases reduce the LLM’s ability to safely respond to harmful contexts residing in completions. Attackers can exploit these biases by inducing the model to generate specific harmful contexts, facilitating a jailbreak attack. Experiments verifying these observations are elaborated in Section 5.2.
@@ 3.1 Dialog Modeling and its Discrepancy
A foundational aspect influencing LLMs’ discrepancy of perceiving query and completion lies in their manner of formatting and modeling the dialog. Open-source LLMs employ dialogue templates to organize user queries and model completion, using distinct tokens to separate them. For instance, LLAMA-2 utilizes the below template to format a dialogue, where query and completion are isolated with “[/INST]”.
Dialog Template of LLAMA-2 [INST] <<SYS>> You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, …(ommitted system prompt) <</SYS>> {{USER QUERY}} [/INST] {{LLM COMPLETION}}
This formatting is not a superficial feature; it enables LLMs trained with such dialogue datasets to inherently differentiate between queries and completions, a critical aspect for effective dialogue modeling. For more templates and special tokens open-source LLMs, please refer to Appendix C.
However, this discrepancy unveils potential vulnerabilities, particularly when paired with biased fine-tuning data. Given the distinction between query and completion, the distribution modeled by the LLM, conditioning on the same context in either the query or the completion, exhibits a disparity:
@@
where $x$ represents the context integrated into the query, $x^{\prime}$ refers to the context residing in the completion, and $y$ is the model’s response. This divergence becomes critical when $\pi_{\Theta}(y|x)$ represents the LLM’s safe response to a hazardous context, and $\pi_{\Theta}(y|x^{\prime})$ diverges from $\pi_{\Theta}(y|x)$. Thus, the discrepancy in dialogue modeling potentially hinders the generalization of safe responses to harmful contexts within completions, thus laying the foundation for potential vulnerabilities.
@@ 3.2 Fine-Tuning and its Safety Biases
The bias in fine-tuning stems from the distinct roles of queries and completions within the objective functions. To understand this bias, it is pertinent to examine the objectives of three predominant fine-tuning methodologies: Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF), and Direct Preference Optimization (DPO).
@@
    $\small\begin{split}&L_{SFT}(\Theta)=\\ &\mathbb{E}_{(x,y)\sim\mathcal{D}_{SFT}}\left[\log\pi_{\Theta}(y_{1}|x)+\sum_{i=1}^{m-1}\log\pi_{\Theta}(y_{i+1}|x,y_{1},...,y_{i})\right]\end{split}$ (2)
    where $x$ is the context containing the prompt, which typically includes the system-generated prompt, and $y$ is the reference answer paired with $x$ in the training set $\mathcal{D}_{SFT}$.
  • Reinforcement Learning from Human Feedback: This technique trains a static reward model $r(x,y)$ using datasets based on human preferences. Then, the LLM, referred to as a policy $\pi_{\Theta}$, undergoes training via reinforcement learning methods, predominantly Proximal Policy Optimization (PPO) [52, 30]. The objective function of PPO is:
    $\small\begin{split}&L_{PPO}(\Theta)=\\ &\mathbb{E}_{x\sim\mathcal{D}_{PPO},y\sim\pi_{\Theta}(y|x)}\left[r(x,y)\right]+\beta\mathbb{D}_{KL}\left[\pi_{\Theta}(y|x)||\pi_{ref}(y|x)\right]\end{split}$ (3)
    Here, $y$ is the completion sampled from the distribution of policy $\pi_{\Theta}(y|x)$, $\pi_{ref}$ is the SFT model, and $\beta$ is a coefficient that penalizes deviations of $\pi_{\Theta}$ from the reference policy $\pi_{ref}$. The policy is initialized with the parameters of $\pi_{ref}$, which establishes a foundation for further optimization.
@@
    $\small\begin{split}&L_{DPO}(\Theta)=\\ &\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{DPO}}\left[\log\sigma(\beta\log\frac{\pi_{\Theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\Theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})\right]\end{split}$ (4)
@@
  • Biased distribution of harmful instruction. In SFT and DPO, harmful instructions seldom appear in completions since its natural to train the model to respond in a harmless way. In RLHF, the policy is less likely to generate harmful instructions because it is based on the SFT model. Consequently, LLMs are less exposed to harmful content in completions than in queries.
@@
    $\left\{\begin{split}&p(y=d,x)>p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{declination}\\ &p(y=d,x)<p(y=d,x^{\prime}),~{}\forall~{}d\in\mathcal{D}_{cooperation}\end{split}\right.$ (5)
    Here, $\mathcal{D}_{declination}$ denotes the set comprising all potential responses that decline harmful content, whereas $\mathcal{D}_{cooperation}$ encompasses all responses that potentially facilitate harmful behaviors. Variable $x$ represents the context in which harmful content is integrated within the query, and $x^{\prime}$ refers to the scenario where harmful content presents in the completion.
Due to the inaccessibility of LLMs’ fine-tuning process, these biases are indirectly verified in Section 5.2 by analyzing the behavior of LLMs after safety fine-tuning.